OpenStudy (anonymous):

Use the logarithmic differentiation to find the derivative of the function y= X^2 times cosx

OpenStudy (anonymous):

\[f(x)=x^2\cos(x)\]?

OpenStudy (anonymous):

or it is \[f(x)=x^{2\cos(x)}\] the first one does not need log diff, the second one does

OpenStudy (anonymous):

Well in my textbook it says to find the derivative of the first one using the log diff

OpenStudy (anonymous):

i would use the product rule \[(fg)'=f'g+g'f\] cant imagine what would be gained using logarithms, you would have to clean up a big mess when you were done

OpenStudy (anonymous):

ok, thanks!!!